# LLM4MSR: An LLM-Enhanced Paradigm for Multi-Scenario Recommendation

### v1

@@ Abstract.
As the demand for more personalized recommendation grows and a dramatic boom in commercial scenarios arises, the study on multi-scenario recommendation (MSR) has attracted much attention, which uses the data from all scenarios to simultaneously improve their recommendation performance. However, existing methods tend to integrate insufficient scenario knowledge and neglect learning personalized cross-scenario preferences, thus leading to sub-optimal performance and inadequate interpretability. Meanwhile, though large language model (LLM) has shown great capability of reasoning and capturing semantic information, the high inference latency and high computation cost of tuning hinder its implementation in industrial recommender systems. To fill these gaps, we propose an effective efficient interpretable LLM-enhanced paradigm LLM4MSR in this work. Specifically, we first leverage LLM to uncover multi-level knowledge including scenario correlations and users’ cross-scenario interests from the designed scenario- and user-level prompt without fine-tuning the LLM, then adopt hierarchical meta networks to generate multi-level meta layers to explicitly improves the scenario-aware and personalized recommendation capability. Our experiments on KuaiSAR-small, KuaiSAR, and Amazon datasets validate two significant advantages of LLM4MSR: (i) the effectiveness and compatibility with different multi-scenario backbone models (achieving 1.5%, 1%, and 40% AUC improvement on three datasets), (ii) high efficiency and deployability on industrial recommender systems, and (iii) improved interpretability. The implemented code and data is available to ease reproduction [^1] https://anonymous.4open.science/r/LLM4MSR/.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information systems Recommender systems
@@ 1. Introduction
Nevertheless, existing MSR methods usually suffer from the following two limitations: (i) Insufficient scenario knowledge is incorporated. Specifically, domain indicator is usually the only scenario knowledge leveraged to capture the domain distinction in practice (Sheng et al., 2021; Chang et al., 2023), while the abundant semantic information such as the description of scenario is neglected. Consequently, the correlations among scenarios are not adequately established. (ii) Users’ personalized preferences across scenarios tend to be ignored, since most MSR models merely rely on different parameter sharing patterns in multi-task learning (Sheng et al., 2021; Wang et al., 2023a) and the collaborative signal learned in conventional RSs to conduct recommendation. However, these deficiencies are still overlooked even if they would lead to the sub-optimal performance of multi-scenario recommender system.
To tackle these drawbacks of conventional MSR models, we resort to large language model (LLM) which is known for its remarkable capability of language understanding and reasoning (Fan et al., 2023). Although there have been some attempts to enhance RSs harnessing LLM (Gong et al., 2023; Yin et al., 2023; Li et al., 2023), there are some key challenges. The first is high inference latency of LLM since the existing methods conduct training and inference with LLM through sample level alignment (Li et al., 2023; Wang et al., 2023b). Second, the information generated from LLM is usually in natural language format. In other words, it requires filling the gap between semantic and recommendation space like employing additional knowledge encoder (Xi et al., 2023) and instruction tuning on LLM (Yin et al., 2023) which brings high computation cost. Moreover, both conventional RS and LLM-based recommendation methods lack interpretability.
@@ 2.1. Multi-Scenario CTR Prediction
[…]
@@
2.2. Multi-Scenario Backbone Models
In the recommendation community, multi-task learning (MTL) technique is first developed targeting at leveraging a unified framework to learn the inter-related tasks like click and conversion, which would lead to mutual enhancement among tasks and higher efficiency (Wang et al., 2023a). Specifically, a mainstream solution is to design various parameter sharing mechanisms with shared and specific parameters in multi-task modeling. Afterward, most multi-scenario recommendation models adopt a similar approach when modeling scenario correlations by simply regarding scenarios as tasks. They model the commonality and distinction among scenarios through explicitly maintaining domain-shared parameters $\Theta_{s}$ and domain-specific parameters $\Theta_{u}$. As depicted in Figure 1, the feature embedding is first fed into the shared parameters like the experts in MMoE (Ma et al., 2018) and the centered network in STAR (Sheng et al., 2021) to extract common knowledge. Then the scenario-specific parameters like gates (Tang et al., 2020) and towers are employed to generate the prediction result.
@@ 2.2. Multi-Scenario Backbone Models
2.3. Large Language Model
Unlike Pre-trained Language Models (PLMs) like BERT trained with masked language modeling task, most popular Large Language Models like GPT (Brown et al., 2020) and ChatGLM2 (Du et al., 2022) adopt the decoder-only architecture and they are trained with next token prediction objective in a self-supervised manner, i.e., a sequence of text is taken as input to predict the next token for each token in the sequence:
(2) $\mathcal{L}_{\theta}=-\sum_{i=1}^{N}\log P_{\theta}\left(\bm{u}_{i}~{}|~{}\bm{u}_{1},\dots,\bm{u}_{i-1}\right)$
where $\bm{u}$ is the input text with $N$ tokens and $\theta$ is the parameters of LLM. The probability distribution of future tokens conditioned on the previous ones are maximized, which is equivalent to minimizing the negative log-likelihood above.
@@ 2.3. Large Language Model
Meanwhile, it adopts masked self-attention or the so-called causal self-attention preventing accessing future tokens, which enables to understand the text data and generate proper results. Specifically, a query (Q), a key (K), and a value (V) vector are obtained multiplying the input token vector $\bm{X}$ by weight matrices $\bm{W}^{Q}$, $\bm{W}^{K}$, and $\bm{W}^{V}$. Then a score is calculated multiplying the current query vector and key vector, which is further multiplied by the value vector and summed up:
(3) $\operatorname{Attenion}(\bm{X})=\bm{W}^{V}\bm{X}\cdot\operatorname{Softmax}\left[\frac{\left(\bm{W}^{K}\bm{X}\right)^{\top}\left(\bm{W}^{Q}\bm{X}\right)}{\sqrt{d_{k}}}+\bm{M}\right]$
where $\bm{M}$ is the mask matrix. In practice, the raw text sequence is passed through the tokenizer to obtain tokens, which are embeded and passed into the decoder-only transformer with several feed-forward and masked self-attention blocks. Then the hidden vectors with the maximum probability are produced and decoded to generate the results in an auto-regressive fashion.
3. Proposed Method
We first illustrate the overall framework of the proposed LLM4MSR paradigm. Next, it is further detailed in the multi-scenario knowledge reasoning and multi-level knowledge fusion step. Finally, we discuss the idea and characteristics of LLM4MSR compared with the existing multi-scenario and LLM-based recommendation methods, and illustrate the necessity of LLMs versus PLMs.
@@ 3.3. Multi-Level Knowledge Fusion
Existing methods either choose to directly leverage the augmented information from LLM as additional input features in the RS (Xi et al., 2023), adopt a linear projection or a fully connected network to conduct dimension transformation and align instance representation (Li et al., 2023), or fine-tune the LLM (Zheng et al., 2023). However, we argue that these are sub-optimal solutions (Su et al., 2023). First, since the LLM-augmented information is treated as a common feature by recommender system, it is unable to play a significant role as expected because the number of feature fields is usually large (Zhao et al., 2021). Second, simply aligning the dimension would result in loss of information (Zebari et al., 2020). Third, tuning LLM brings high cost (Chen et al., 2023b, a) and it requires consistent update.
@@
where $\bm{W}_{l}^{(i)}$ and $\bm{b}_{l}^{(i)}$ indicate the weight and bias matrix of the $i$-th meta layer. Notably, $Dim_{mw}$ and $Dim_{mb}$ equals to the sum of dimensions of all $\bm{W}_{l}^{(i)}$ and $\bm{b}_{l}^{(i)}$, respectively. For example, $\bm{h}_{mw}$ with the dimension of 8256 can be reshaped into a weight matrix with dimension 128$\times$64 plus a weight matrix with dimension 64$\times$1 (bias matrix is similar). Then each meta layer is followed by an activation function $\sigma$ like ReLU:
@@ 3.4. Discussion
LLM4MSR v.s. LLM-based RS. Existing LLM-based recommender systems can be categorized into two groups, i.e., using LLM as the knowledge enhancer (Li et al., 2023; Wang et al., 2023b; Lin et al., 2023; Gong et al., 2023) and directly as the recommender (Geng et al., 2022; Cui et al., 2022; Liu et al., 2023). However, the former models possess low efficiency since they require aligning the information from LLM and RS in the instance level. Meanwhile, the latter methods lack scenario knowledge, thus achieving unsatisfying recommendation performance and suffering from high inference latency of LLM.
By contrast, LLM4MSR can efficiently and effectively tackle these problems. First, thanks to the zero-shot capability (Wang et al., 2022b), an LLM with fixed parameters is leveraged free from tuning in (Yin et al., 2023; Li et al., 2023; Wang et al., 2023b). Meanwhile, the way of obtaining information from LLM is shifted from the low ‘instance-level’ (e.g., in (Li et al., 2023; Wang et al., 2023b; Zheng et al., 2023)) to the high ‘user- & scenario-level’, thus improving efficiency. Besides, through hierarchical meta networks, the multi-level scenario knowledge is adaptively injected and lead to enhancement. Moreover, apart from distinct scenario-specific and user-specific features, it also enables to incorporate expert scenario knowledge through the prompt design, which is more universal and advanced than simply extracting invariant features and aligning the heterogeneous ones in (Gong et al., 2023).
Necessity of LLMs v.s. PLMs. As illustrated in Section 2.3, Different training objectives of Large Language Models (LLMs) and Pre-trained Language Models (PLMs) determine their roles. Specifically, rather than simply deriving a representation of existing (old) information by adopting PLMs as encoder (like BERT in CTRL (Li et al., 2023)), the key of leveraging LLM as reasoner in LLM4MSR is to further infer useful (new) knowledge like analyzing user’s cross-scenario preference. This knowledge (in the form of hidden vectors) could potentially help enhancing conventional recommender systems. Besides, a case study is conducted in Section 4.6 and the knowledge (in the form of text) from ChatGLM2-6B is shown in Figure 5, which helps improving interpretability.
@@ 4.1.1. Datasets
Scenarios Rec Search Rec Search Rec#$\#1$ Rec#$\#2$ Rec#$\#3$
Users 25877 25877 25877 25877 24752 24752 24752
Items 2281034 1974165 4046363 2974596 8788 193304 6980
Interactions 7493101 3038362 14605712 4828690 33929 370840 56356
@@ 4.1.3. Backbone Models
Table 2. Overall performance comparison of AUC on KuaiSAR-small, KuaiSAR, and Amazon dataset. The suffix ‘DN’, ‘EP’, and ‘ours’ denote Dynamic Network, EPNet, and LLM4MSR. Boldface denotes the highest value while underline indicates the second best result. $\star$ represents statistical significance with $p$-value $<0.05$ in $t$-test compared with the backbone model.
@@ 4.4. Hyper-parameter Analysis (RQ3)
We analyze the effect of two significant hyper-parameters, i.e., the threshold ($T$) of positive interactions in the user-level prompt and the number of fully connected layers ($K$) generated by the meta network using STAR as the backbone model on KuaiSAR-small dataset. Specifically, as shown in Figure 4(a), the recommendation accuracy continuously increases as $T$ rises until the token limit of LLM is reached, since more information about cross-scenario preference of users is introduced and captured by LLM. Besides, from Figure 4(b) we can see that the best performance is obtained where $K=2$. Nevertheless, the results degrade as the network becomes deeper and a possible reason is overfitting.
@@ 4.6. Case Study (RQ5)
Table 3. Efficiency comparison on KuaiSAR-small in seconds.
Models STAR STAR_DN STAR_EP STAR_ours
Training 2.8e-2 2.9e-2 2.9e-2 1.5e-1
Inference 1.7e-2 1.7e-2 1.7e-2 5.2e-2
@@ 5.1. Multi-Scenario Recommendation
Multi-scenario recommendation (Wang et al., 2023c, 2022a), also known as multi-domain or multi-target cross-domain recommendation (Zhu et al., 2021), uses the data from all scenarios to simultaneously improve their recommendation performance and tackle data sparsity problem. For example, STAR (Sheng et al., 2021) adopts a star-like framework with a shared centered network and scenario-specific networks similar to the parameter sharing pattern design in multi-task modeling (Wang et al., 2023a). M2M (Zhang et al., 2022) proposes a meta learning mechanism with hierarchically organized meta attention and meta tower module to grasp inter-scenario correlations and scenario-specific representation. In addition, to address the imperfectly double seesaw phenomenon, PEPNet (Chang et al., 2023) proposes embedding and parameter personalized network to dynamically scale the parameters based on the designed gate mechanism in a domain-specific and task-specific manner, respectively.
@@ 5.2. Language Model for CTR Prediction
Several works try to incorporate Language Models (LMs) including Large Language Models (LLMs) and Pre-trained Language Models (PLMs) as illustrated in Section 2.3 with recommendation for Click-Through Rate (CTR) prediction (Li et al., 2023; Xi et al., 2023; Lin et al., 2023; Wang et al., 2023b; Gong et al., 2023) since they have the potential to bring information gain. Specifically, one of the solutions is to treat the semantic knowledge as cross-modal information and align with the collaborative signal in conventional recommender systems. For example, CTRL (Li et al., 2023) employs contrastive learning for knowledge alignment and integration. KAR (Xi et al., 2023) first generates reasoning knowledge of user preferences and item factual information, then encodes and adapts them to the backbone recommendation models. Besides, S&R Multi-domain Foundation Model (Gong et al., 2023) simply adopts LLM to extract scenario-invariant text features and tackle item heterogeneity issue.

### v2

@@ Abstract.
As the demand for more personalized recommendation grows and a dramatic boom in commercial scenarios arises, the study on multi-scenario recommendation (MSR) has attracted much attention, which uses the data from all scenarios to simultaneously improve their recommendation performance. However, existing methods tend to integrate insufficient scenario knowledge and neglect learning personalized cross-scenario preferences, thus leading to sub-optimal performance. Meanwhile, though large language model (LLM) has shown great capability of reasoning and capturing semantic information, the high inference latency and high computation cost of tuning hinder its implementation in industrial recommender systems. To fill these gaps, we propose an LLM-enhanced paradigm LLM4MSR in this work. Specifically, we first leverage LLM to uncover multi-level knowledge from the designed scenario- and user-level prompt without fine-tuning the LLM, then adopt hierarchical meta networks to generate multi-level meta layers to explicitly improve the scenario-aware and personalized recommendation capability. Our experiments on KuaiSAR-small, KuaiSAR, and Amazon datasets validate significant advantages of LLM4MSR: (i) the effectiveness and compatibility with different multi-scenario backbone models, (ii) high efficiency and deployability on industrial recommender systems, and (iii) improved interpretability. The implemented code and data is available to ease reproduction [^1] https://github.com/mindspore-lab/models/tree/master/research/huawei-noah/LLM4MSR, [^2] https://github.com/Applied-Machine-Learning-Lab/LLM4MSR.
@@
🖂Corresponding author
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679743isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systems
@@ 1. Introduction
Nevertheless, existing MSR methods usually suffer from the following two limitations: (i) Insufficient scenario knowledge is incorporated. Specifically, domain indicator is usually the only scenario knowledge leveraged to capture the domain distinction in practice (Sheng et al., 2021; Chang et al., 2023), while the abundant semantic information such as the description of scenario is neglected. Consequently, the correlations among scenarios are not adequately established. (ii) Users’ personalized preferences across scenarios tend to be ignored, since most MSR models merely rely on different parameter sharing patterns in multi-task learning (Sheng et al., 2021; Wang et al., 2023a; Liu et al., 2023b) and the collaborative signal learned in conventional RSs to conduct recommendation. However, these deficiencies are still overlooked even if they would lead to the sub-optimal performance of multi-scenario recommender system.
To tackle these drawbacks of conventional MSR models, we resort to large language model (LLM) which is known for its remarkable capability of language understanding and reasoning (Fan et al., 2023). Although there have been some attempts to enhance RSs harnessing LLM (Gong et al., 2023; Yin et al., 2023; Li et al., 2023a), there are some key challenges. The first is high inference latency of LLM since the existing methods conduct training and inference with LLM through sample level alignment (Li et al., 2023a; Wang et al., 2023b). Second, the information generated from LLM is usually in natural language format. In other words, it requires filling the gap between semantic and recommendation space like employing additional knowledge encoder (Xi et al., 2023) and instruction tuning on LLM (Yin et al., 2023) which brings high computation cost. Moreover, both conventional RS and LLM-based recommendation methods lack interpretability.
@@ 2.1. Multi-Scenario CTR Prediction
The Click-Trough Rate (CTR) prediction task is essential in recommendation which is a binary classification problem. Specifically, in a multi-scenario recommender system, the historical interaction sample $(d,\bm{x})$ is taken as input to predict the ground truth label $y$ where $y=1$ and $y=0$ denote click and unclick, respectively. $d$ represents the domain indicator $d\in\{1,2,\dots,D\}$ indicating the origin of sample. $\bm{x}$ denotes the feature fields including user and item attributes. For simplicity, suppose there are $M$ categorical features, then $\bm{x}=[\bm{x}_{1},\ldots,\bm{x}_{m},\ldots,\bm{x}_{M}]$ where $\bm{x}_{m}$ indicates the one-hot encoding of the $m$-th feature. Next, $\bm{x}$ is mapped into a dense vector $\bm{e}=[\bm{e}_{1}\|\ldots\|\bm{e}_{m}\|\ldots\|\bm{e}_{M}]$ through an embedding layer where $\|$ denotes the concatenation operation. For the $m$-th feature, we calculate $\bm{e}_{m}$ through a look-up operation $\bm{e}_{m}=\bm{E}_{m}\cdot\bm{x}_{m}$ in which $\bm{E}_{m}\in\mathbb{R}^{u_{m}\times Dim}$ denotes the embedding table, $u_{m}$ denotes the value counts of feature, and $Dim$ is the embedding size. Finally, the prediction result $\hat{y}$ is obtained via $\hat{y}=f_{d}(\bm{e})$ where $f_{d}$ represents the recommendation model applied to the $d$-th scenario. The loss function is binary cross entropy loss or the so-called Logloss (Berkson, 1944):
@@
As depicted in Figure 1, most multi-scenario recommendation models model the commonality and distinction among scenarios through explicitly maintaining domain-shared parameters $\Theta_{s}$ like the experts in MMoE (Ma et al., 2018) and domain-specific parameters $\Theta_{u}$ like towers and gates in PLE (Tang et al., 2020).
@@
3. Proposed Method
@@
We first illustrate the overall framework of LLM4MSR. Next, it is further detailed in the multi-scenario knowledge reasoning and multi-level knowledge fusion step. Finally, we discuss the idea and characteristics of LLM4MSR compared with the existing multi-scenario and LLM-based recommendation methods, and illustrate the necessity of LLMs versus Pre-trained Language Models (PLMs).
@@ 3.3. Multi-Level Knowledge Fusion
Existing methods either choose to directly leverage the augmented information from LLM as additional input features in the RS (Xi et al., 2023), adopt a linear projection or a fully connected network to conduct dimension transformation and align instance representation (Li et al., 2023a), or fine-tune the LLM (Zheng et al., 2023). However, we argue that these are sub-optimal solutions (Su et al., 2023). First, since the LLM-augmented information is treated as a common feature by recommender system, it is unable to play a significant role as expected because the number of feature fields is usually large (Zhao et al., 2021). Second, simply aligning the dimension would result in loss of information (Zebari et al., 2020). Third, tuning LLM brings high cost (Chen et al., 2023b, a) and it requires consistent update.
@@
[…]
@@ 3.4. Discussion
LLM4MSR v.s. LLM-based RS. Existing LLM-based recommender systems can be categorized into two groups, i.e., using LLM as the knowledge enhancer (Li et al., 2023a; Wang et al., 2023b; Lin et al., 2023; Gong et al., 2023) and directly as the recommender (Geng et al., 2022; Cui et al., 2022; Liu et al., 2023a). However, the former models possess low efficiency since they require aligning the information from LLM and RS in the instance level. Meanwhile, the latter methods lack scenario knowledge, thus achieving unsatisfying recommendation performance and suffering from high inference latency of LLM.
By contrast, LLM4MSR can efficiently and effectively tackle these problems. First, thanks to the zero-shot capability (Wang et al., 2022b), an LLM with fixed parameters is leveraged free from tuning in (Yin et al., 2023; Li et al., 2023a; Wang et al., 2023b). Meanwhile, the way of obtaining information from LLM is shifted from the low ‘instance-level’ (e.g., in (Li et al., 2023a; Wang et al., 2023b; Zheng et al., 2023)) to the high ‘user- & scenario-level’, thus improving efficiency. Besides, through hierarchical meta networks, the multi-level scenario knowledge is adaptively injected and lead to enhancement. Moreover, apart from distinct scenario-specific and user-specific features, it also enables to incorporate expert scenario knowledge through the prompt design, which is more universal and advanced than simply extracting invariant features and aligning the heterogeneous ones in (Gong et al., 2023).
Necessity of LLMs v.s. PLMs. Unlike PLMs like BERT trained with masked language modeling task, most popular LLMs like ChatGLM2 (Du et al., 2022) are trained with next token prediction objective. Therefore, rather than simply deriving a representation of existing (old) information by adopting PLMs as encoder (like BERT in CTRL (Li et al., 2023a)), the key of leveraging LLM as reasoner in LLM4MSR is to further infer useful (new) knowledge like analyzing user’s cross-scenario preference. This knowledge (in the form of hidden vectors) could potentially help enhancing conventional recommender systems. Besides, a case study is conducted in Section 4.6 and the knowledge (in the form of text) from ChatGLM2-6B is shown in Figure 5, which helps improving interpretability.
@@ 4.1.1. Datasets
Scenarios Rec Search Rec Search Rec#1\#1 Rec#2\#2 Rec#3\#3
Users 25,877 25,877 25,877 25,877 24,752 24,752 24,752
Items 2,281,034 1,974,165 4,046,363 2,974,596 8,788 193,304 6,980
Interactions 7,493,101 3,038,362 14,605,712 4,828,690 33,929 370,840 56,356
@@ 4.1.3. Backbone Models
Table 2. Overall performance comparison of AUC on KuaiSAR-small, KuaiSAR, and Amazon dataset. The suffix ‘DN’, ‘EP’, and ‘ours’ denote Dynamic Network, EPNet, and LLM4MSR. Boldface denotes the highest value while underline indicates the second best result. $\star$ represents statistical significance with $p$-value <0.05<0.05 in $t$-test compared with the backbone model.
@@ 4.4. Hyper-parameter Analysis (RQ3)
We analyze the effect of two significant hyper-parameters, i.e., the threshold ($T$) of positive interactions in the user-level prompt and the number of fully connected layers ($K$) generated by the meta network using STAR as the backbone model on KuaiSAR-small dataset. Specifically, as shown in Figure 4(a), the recommendation accuracy continuously increases as $T$ rises until the token limit of LLM is reached, since more information about cross-scenario preference of users is introduced and captured by LLM. Besides, from Figure 4(b) we can see that the best performance is obtained where $K=2$. Nevertheless, the results degrade as the network becomes deeper and a possible reason is over-fitting.
@@ 4.5. Efficiency Analysis (RQ4)
Table 3. Efficiency comparison on KuaiSAR-small in seconds.
Models STAR STAR_DN STAR_EP STAR_ours
Training 2.8e-2 2.9e-2 2.9e-2 1.5e-1
Inference 1.7e-2 1.7e-2 1.7e-2 5.2e-2
@@ 5.1. Multi-Scenario Recommendation
Multi-scenario recommendation (Wang et al., 2023c, 2022a, 2024; Li et al., 2023c), also known as multi-domain or multi-target cross-domain recommendation (Zhu et al., 2021), uses the data from all scenarios to simultaneously improve their recommendation performance and tackle data sparsity problem. For example, STAR (Sheng et al., 2021) adopts a star-like framework with a shared centered network and scenario-specific networks similar to the parameter sharing pattern design in multi-task modeling (Wang et al., 2023a). M2M (Zhang et al., 2022) proposes a meta learning mechanism with hierarchically organized meta attention and meta tower module to grasp inter-scenario correlations and scenario-specific representation. In addition, to address the imperfectly double seesaw phenomenon, PEPNet (Chang et al., 2023) proposes embedding and parameter personalized network to dynamically scale the parameters based on the designed gate mechanism in a domain-specific and task-specific manner, respectively.
@@ 5.2. Language Model for CTR Prediction
Several works try to incorporate Language Models (LMs) including Large Language Models (LLMs) and Pre-trained Language Models (PLMs) with recommendation for Click-Through Rate (CTR) prediction (Li et al., 2023a; Fu et al., 2023; Xi et al., 2023; Lin et al., 2023; Wang et al., 2023b; Gong et al., 2023) since they have the potential to bring information gain (Zhao et al., 2024; Xu et al., 2023; Li et al., 2023b; Liu et al., 2024). Specifically, one of the solutions is to treat the semantic knowledge as cross-modal information and align with the collaborative signal in conventional recommender systems. For example, CTRL (Li et al., 2023a) employs contrastive learning for knowledge alignment and integration. KAR (Xi et al., 2023) first generates reasoning knowledge of user preferences and item factual information, then encodes and adapts them to the backbone recommendation models. Besides, S&R Multi-domain Foundation Model (Gong et al., 2023) simply adopts LLM to extract scenario-invariant text features and tackle item heterogeneity issue.